OpenStudy (cutiekawdfi):

solve by factoring x^2 + 2x = 8

OpenStudy (anonymous):

subtract 8 from both sides x^2 + 2x - 8=0 (x-a)(x-b)=0 ab=-8 and a+b=-2

OpenStudy (anonymous):

if a = -4 and b=2, the equations work

OpenStudy (anonymous):

(x-(-4))(x-2)=0 (x+4)(x-2)=0 solutions are -4 and 2

OpenStudy (cutiekawdfi):

thank youu